OpenStudy (anonymous):

Find the area of a regular pentagon with a side length of 14 and an apothem of 22.

OpenStudy (tkhunny):

360/5 = 72 72/2 = 36 180 - 72 = 108 108/2 = 54 14/2 = 7  Just draw it and do the arithmetic. How many triangles like that are there in a Regular Pentagon?